# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v2

@@ Abstract
The use of copyrighted books for training AI models has led to numerous lawsuits from authors concerned about AI’s ability to generate derivative content. Yet it’s unclear whether these models can generate high quality literary text while emulating authors’ styles/voices. To answer this we conducted a preregistered study comparing MFA-trained expert writers with three frontier AI models: ChatGPT, Claude, and Gemini in writing up to 450 word excerpts emulating 50 award-winning authors’ (including Nobel laureates, Booker Prize winners, and young emerging National Book Award finalists) diverse styles. In blind pairwise evaluations by 159 representative expert (MFA candidates from top U.S. writing programs) and lay readers (recruited via Prolific), AI-generated text from in-context prompting was strongly disfavored by experts for both stylistic fidelity (odds ratio [OR]=0.16, $p<10^{-8}$) and writing quality (OR=0.13, $p<10^{-7}$) but showed mixed results with lay readers. However, fine-tuning ChatGPT on individual author’s complete works completely reversed these findings: experts now favored AI-generated text for stylistic fidelity (OR=8.16, $p<10^{-13}$) and writing quality (OR=1.87, p=0.010), with lay readers showing similar shifts. These effects are robust under cluster-robust inference and generalize across authors and styles in author-level heterogeneity analyses. The fine-tuned outputs were rarely flagged as AI-generated (3% rate versus 97% for in-context prompting) by state-of-the-art AI detectors. Mediation analysis reveals this reversal occurs because fine-tuning eliminates detectable AI stylistic quirks (e.g., cliché density) that penalize in-context outputs, altering the relationship between AI detectability and reader preference. While we do not account for additional costs of human effort required to transform raw AI output into cohesive, publishable prose, the median fine-tuning and inference cost of \$81 per author represents a dramatic 99.7% reduction compared to typical professional writer compensation. Author-specific fine-tuning thus enables non-verbatim AI writing that readers prefer to expert human writing, thereby providing empirical evidence directly relevant to copyright’s fourth fair-use factor, the “effect upon the potential market or value” of the source works.
@@
1 Introduction
The U.S. publishing industry supports hundreds of thousands of jobs while generating \$30 billion in yearly revenue, contributing to the larger American copyright sectors that account for \$2.09 trillion in annual GDP contributions (?). Adult Fiction and nonfiction books alone accounted for \$6.14 billion in 2024 (?). This economically important sector now faces an unprecedented challenge: its core products have become essential training data for generative-AI systems. Models trained on well-edited books produce more coherent, accurate responses—something crucial to creating the illusion of intelligence (?). Most technology companies building AI use massive datasets of books, typically without permission or licensing (?), and frequently from illegal sources. In the recent copyright lawsuit Bartz v. Anthropic (?), Judge Alsup noted that Anthropic acquired at least five million books from LibGen and two million from Pirate Library Mirror (PiLiMi). Anthropic also used Books3111Now removed after a legal complaint by anti-piracy group, the Rights Alliance.—a dataset of approximately 191,000 books also used by Meta and Bloomberg to train their language models. [^2] The Bartz v. Anthropic lawsuit also revealed how Anthropic cut millions of print books from their bindings, scanned them into digital files, and threw away the originals solely for the purpose of training Claude. This unauthorized use has sparked outrage among authors (?), triggering dozens of lawsuits against technology companies including OpenAI, Anthropic, Microsoft, Google, and Meta.
@@ 1 Introduction
While it is widely established that most frontier large language models (LLMs) have been trained on copyrighted books, it remains unclear whether such training can produce expert-level creative writing. Past research has shown that AI cannot produce highbrow literary fiction or creative nonfiction through prompting alone when compared to professionally trained writers (?). More recent work from (?) demonstrates that AI-generated creative writing still remains characterized by clichés, purple prose, and unnecessary exposition. Additionally, relying on GenAI for creative writing reduces the collective diversity of novel content (?). AI often produces formulaic, mediocre creative writing because it lacks the distinctive personal voice that typically distinguishes one author from another (?). As Pulitzer Fiction finalist Vauhini Vara observes, “ChatGPT’s voice is polite, predictable, inoffensive, upbeat. Great characters, on the other hand, aren’t polite; great plots aren’t predictable; great style isn’t inoffensive; and great endings aren’t upbeat” (?). To address this limitation, practitioners now increasingly prompt AI systems to perform style/voice mimicry by emulating specific writers’ choices (?). This practice has become so common that a fantasy author recently published a novel containing an accidentally included AI prompt requesting emulation of another writer’s style (?). While the effectiveness of such stylistic emulation remains contested, the more pressing question concerns whether style/voice mimicry genuinely improves AI-generated text quality and whether readers—both experts and non-experts—perceive these improvements as meaningful.
To address this question, we conducted a preregistered behavioral study comparing MFA-trained expert writers with froniter large language models. Historically, top MFA programs have produced many prizewinning American writers  (?). Eminent literary agent Gail Hochman of Brandt & Hochman said, “We look favorably on anyone who has an M.F.A., simply because it shows they’re serious about their writing”  (?). This elite sample of MFA trained expert writers provides a conservative test—if AI can compete with the best emerging talent, the disruption to average writers is likely even greater. We selected closed-source LLMs readily accessible to users without technical expertise: GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. [^3] We also tried open-weight Llama 3.1 model but its empirical performance was not good at following long context instructions at the time of our study. Additionally our results would hold across different LLM versions in the future as LLM outputs have been shown to be homogeneous and have not improved over time (?, ?, ?, ?). Both human experts and LLMs were given the same task: write an excerpt of up to 450 words emulating the style and voice of one of 50 internationally acclaimed authors, including Nobel laureates, Booker Prize winners, and Pulitzer Prize winners, spanning multiple continents and cultures. [^4] Among others, our study includes Nobel laureates Han Kang and Annie Ernaux; Booker Prize winners Salman Rushdie, Margaret Atwood, and George Saunders; and Pulitzer Prize winners Junot Diaz and Marilynne Robinson. We tested two AI conditions: (1) in-context prompting, where models received the same instructions as human experts, and (2) fine-tuning, where models were additionally trained on each author’s complete oeuvre. [^5] For authors writing in non-English languages (Han Kang, Yoko Ogawa, Annie Ernaux, Haruki Murakami), we used the same translator’s work across all books to maintain voice consistency. Expert and lay readers performed blind pairwise evaluations (?, ?, ?) of <<Human-AI>> excerpts on writing quality and stylistic fidelity. This design addresses three preregistered research questions: (1) Can AI match expert performance in writing quality and stylistic fidelity across both conditions? (2) Do expert and lay readers show similar preference patterns? (3) Does AI detectability correlate with human quality judgments, and does fine-tuning remove this correlation? Our full experimental setup is shown in Figure 1.
@@ 2 Results
Figure 2: (A-B) Forest plots showing odds ratios (OR) and 95% confidence intervals comparing AI and human experts in pairwise evaluations of stylistic fidelity (A) and writing quality (B) where values >>1 favor AI and values >>1 favor humans. Expert readers show preference for human writing when prompted in an in-context setting (OR = 0.16 and 0.13) but that changes when AI is fine-tuned (OR = 8.16 and 1.87). Lay readers have a harder time discriminating, given how they prefer the quality of AI writing even for in-context prompting (OR = 1.55). (C-D) Probability of choosing AI excerpts across individual language models for stylistic fidelity (C) and writing quality (D). Error bars represent 95% confidence intervals. Dashed line indicates chance performance (50%). (E) AI detection accuracy with chosen threshold of $\tau$=0.9 using two state-of-the-art AI detectors (Pangram and GPTZero). Human written text was never misclassified (0.00), in-context AI was detected with 97% accuracy by Pangram and 91% by GPTZero, but fine-tuned AI evaded detection 97% of the time (0.03) for Pangram and 100% of the time in case of GPTZero. (F) Relationship between AI detectability (Pangram) and preference for writing quality across detection score bins. For in-context prompting setup, higher detection scores correlated with lower AI preference (negative slope). This relationship disappeared when AI is fine-tuned (flat slopes). Fine-tuning on an author’s complete oeuvre eliminates stylometric “AI” quirks while achieving expert-level performance. n = 28 expert readers, 131 lay readers; 3,840 pairwise comparisons with robust clustered standard errors.
@@ 2.1 Overall Performance Comparisons
Our final data consist of 3,840 paired-choice tasks, with judgments from 28 MFA experts and 131 lay readers. We fit a logit model for each outcome and condition and include dummies for writer type, reader group, and their interaction. We further employ CR2 cluster-robust standard errors (?) clustered at the reader-level to account for within-reader correlation in ratings. Our hypotheses, outcomes, design, and analysis closely follow our OSF pre-registration (SI Sections S4-S8); deviations are detailed in SI Section S9.
Figure 2A–B presents the odds ratios, with corresponding predicted probabilities shown in Figure 2C–D. Under in-context prompting, expert readers demonstrated strong preference for human-written text. Odds ratios were 0.16 (95% CI: 0.08–0.29, $p<10^{-8}$) for stylistic fidelity and 0.13 (95% CI: 0.06–0.28, $p<10^{-7}$) for writing quality, indicating six- to eight-fold preferences for human excerpts (Fig. 2A–B). Lay readers showed no significant preference regarding stylistic fidelity (OR = 0.86, 95% CI: 0.62–1.19, $p=0.37$) but favored AI-generated text for writing quality (OR = 1.55, 95% CI: 1.09–2.21, $p=0.014$), selecting AI excerpts in 61% of writing quality trials (Fig. 2C–D). Inter-rater agreement reflected this divergence: expert readers achieved $\kappa=0.58$ for stylistic fidelity and $\kappa=0.41$ for writing quality, while lay readers showed minimal agreement among themselves ($\kappa=0.12$ and $\kappa=0.15$, respectively). [^6] Given that lay assessments of literary quality and style reflect inherently diverse tastes, we anticipated lower inter-rater agreement compared to expert readers. The writer-type $\times$ reader-type interaction was significant for both outcomes ($\chi^{2}_{(3)}=24.9$, $p=1.6\times 10^{-5}$ for fidelity; $\chi^{2}_{(3)}=37.6$, $p=3.5\times 10^{-8}$ for quality).
@@ 2.2 AI Detection and Stylometric Analysis
We probe whether differences in AI detectability can account for these preference reversals. Pangram, a state-of-the-art AI detection tool (?, ?, ?), correctly classified 97% of in-context prompted texts as machine-generated but only 3% of fine-tuned texts were classified as AI-generated (Fig. 2E). [^7] GPTZero, another state-of-the-art AI detection tool showed comparable performance but had higher false-positive-rate, so for our subsequent analyses we stuck to Pangram.
@@ 2.3 Author-Level Performance Heterogeneity
Figure 4: Fine-tuning substantially reduces stylometric signatures of AI text, improves stylistic fidelity and perceived writing quality over in-context prompting, and substantially cuts costs of producing first draft versus professional writers. (A) Mediation analysis linking stylometric features $\rightarrow$ AI-detector score $\rightarrow$ human preference. Standardized logistic coefficients with 95% CIs are shown for three features for in-context prompting (red) and fine-tuned models (green). Cliché density mediates 16.4% of the detector effect on choice for in-context prompting but only 1.3% for author fine-tuned models; all three features together mediate 25.4% vs -3.2% for in-context prompted and fine-tuned models respectively. (B) “Fine-tuning premium,” defined as $\Delta$ = P(prefer fine-tuned over human) - P(prefer in-context over human), as a function of fine-tuning corpus size. Top: stylistic fidelity; bottom: writing quality. Points are authors; colors denote improvement (green, $\Delta$¿0), no change (gray), or degradation (red, $\Delta<0$). Median $\Delta$: +41.7 (fidelity) and +16.7 percentage points (quality). (C) Cost to produce 100,000 words of raw text vs. publishable prose. Expert writers in our study would earn \$25,000 for a 100k-word novel-length manuscript (red). By contrast, AI pipelines can generate 100k words of raw text for \$25–\$276 depending on fine-tuning corpus size (green bars = fine-tuning; hatched = in-context prompting, \$3). This figure reflects direct compute/API costs only, not the additional human steering, chunking, and editing required to turn raw AI text into a cohesive publishable work. Authors ordered by total AI cost.
@@ 2.4 Cost Analysis
The weak dependence of performance on fine-tuning corpus size has direct economic implications. Model fine-tuning and inference costs ranged from \$25 to \$276 per author (median = \$81), assuming API-based fine-tuning at \$25 per million tokens plus \$3 for generating 100,000 words of raw text (Fig. 4C). These costs represent approximately 0.3% of what expert writers in our study would charge for a novel-length (100,000 words) manuscript. It should be noted that this comparison reflects raw generation costs before the human steering and editing required to transform AI outputs into publishable works. Despite this caveat, the minimal investment yielded outputs that achieved expert-level performance for most authors. Performance gains were uncorrelated with both corpus size and fine-tuning cost, indicating that computational scale did not drive improvements. The 99.7% reduction in raw generation costs, coupled with superior quality ratings for the majority of authors, underscores the potential for substantial producer surplus shifts and market displacement.

### v3

@@ Abstract
The use of copyrighted books for training AI models has led to numerous lawsuits from authors concerned about AI’s ability to generate derivative content. Yet it’s unclear whether these models can generate high quality literary text while emulating authors’ styles/voices. To answer this we conducted a preregistered study comparing MFA-trained expert writers with three frontier AI models: ChatGPT, Claude, and Gemini in writing up to 450 word excerpts emulating 50 award-winning authors’ (including Nobel laureates, Booker Prize winners, and young emerging National Book Award finalists) diverse styles. In blind pairwise evaluations by 159 representative expert (MFA candidates from top U.S. writing programs) and lay readers (recruited via Prolific), AI-generated text from in-context prompting was strongly disfavored by experts for both stylistic fidelity (odds ratio [OR]=0.16, $p<10^{-8}$) and writing quality (OR=0.13, $p<10^{-7}$) but showed mixed results with lay readers. However, fine-tuning ChatGPT on individual author’s complete works completely reversed these findings: experts now favored AI-generated text for stylistic fidelity (OR=8.16, $p<10^{-13}$) and writing quality (OR=1.87, p=0.010), with lay readers showing similar shifts. These effects are robust under cluster-robust inference and generalize across authors and styles in author-level heterogeneity analyses. The fine-tuned outputs were rarely flagged as AI-generated (3% rate versus 97% for in-context prompting) by state-of-the-art AI detectors. Mediation analysis reveals this reversal occurs because fine-tuning eliminates detectable AI stylistic quirks (e.g., cliché density) that penalize in-context outputs, altering the relationship between AI detectability and reader preference. While we do not account for additional costs of human effort required to transform raw AI output into cohesive, publishable novel length prose, the median fine-tuning and inference cost of \$81 per author represents a dramatic 99.7% reduction compared to typical professional writer compensation. Author-specific fine-tuning thus enables non-verbatim AI writing that readers prefer to expert human writing, thereby providing empirical evidence directly relevant to copyright’s fourth fair-use factor, the “effect upon the potential market or value” of the source works.
@@
The U.S. publishing industry supports hundreds of thousands of jobs while generating \$30 billion in yearly revenue, contributing to the larger American copyright sectors that account for \$2.09 trillion in annual GDP contributions (?). Adult Fiction and Non-Fiction books alone accounted for \$6.14 billion in 2024 (?). This economically important sector now faces an unprecedented challenge: its core products have become essential training data for generative-AI systems. Models trained on well-edited books produce more coherent, accurate responses—something crucial to creating the illusion of intelligence (?). Most technology companies building AI use massive datasets of books, typically without permission or licensing (?), and frequently from illegal sources. In the recent copyright lawsuit Bartz vs Anthropic (?), Judge Alsup noted that Anthropic acquired at least five million books from LibGen and two million from Pirate Library Mirror (PiLiMi). Anthropic also used Books3 (a dataset of approximately 191,000 books also used by Meta and Bloomberg to train their language models), now removed after a legal complaint by anti-piracy group, the Rights Alliance. The Bartz vs Anthropic lawsuit also revealed how Anthropic cut millions of print books from their bindings, scanned them into digital files, and threw away the originals solely for the purpose of training Claude. This unauthorized use has sparked outrage among authors (?), triggering dozens of lawsuits against technology companies including OpenAI, Anthropic, Microsoft, Google, and Meta.
@@
While it is widely established that most state-of-the-art Large Language Models (LLMs) have been trained on copyrighted books, it remains unclear whether such training can produce expert-level creative writing. Past research has shown that AI cannot produce high-brow literary fiction or creative nonfiction through prompting alone when compared to professionally trained writers (?). More recent work demonstrates that AI-generated creative writing still remains characterized by clichés, purple prose, and unnecessary exposition  (?). Additionally, relying on Generative AI for creative writing reduces the collective diversity of novel content (?). AI often produces formulaic, mediocre creative writing because it lacks the distinctive personal voice that typically distinguishes one author from another (?). As Pulitzer Fiction finalist Vauhini Vara observes, “ChatGPT’s voice is polite, predictable, inoffensive, upbeat. Great characters, on the other hand, aren’t polite; great plots aren’t predictable; great style isn’t inoffensive; and great endings aren’t upbeat” (?). To address this limitation, practitioners now increasingly prompt AI systems to perform style/voice mimicry by emulating specific writers’ choices (?). This practice has become so common that a fantasy author recently published a novel containing an accidentally included AI prompt requesting emulation of another writer’s style (?). While the effectiveness of such stylistic emulation remains contested, the more pressing question concerns whether style/voice mimicry genuinely improves AI-generated text quality and whether readers—both experts and non-experts—perceive these improvements as meaningful.
To address this question, we conducted a preregistered behavioral study comparing a representative sample of MFA-trained expert writers with state-of-the-art LLMs. [^1] In the United States, a Master of Fine Arts (MFA) is typically recognized as a terminal degree for practitioners of creative writing (equivalent to a Ph.D.)—meaning that it is considered the highest degree in its field, qualifying an individual to become a professor at the university level in these disciplines. MFA-trained writers are broadly representative of the universe of expert writers. Not all professional writers attended MFA programs, but enough did to stand in for the wider group. Historically, several award-winning authors who have written great American Novels  (?) have graduated with an MFA degree in Creative Writing (See Figure 9 in SI) and especially from the programs we recruited our writers for this study (See Figure 11 in SI). Eminent literary agent Gail Hochman of Brandt & Hochman said, “We look favorably on anyone who has an M.F.A., simply because it shows they’re serious about their writing”  (?). In June 2010, the editors of The New Yorker Magazine announced to widespread media coverage their selection of 20 under 40222https://www.newyorker.com/magazine/20-under-40-fiction—the young fiction writers who are, or will be, central to their generation; 17 out of those 20 writers had an MFA degree (See Fig 11 in SI). This elite sample of MFA trained expert writers provides a conservative test—if AI can compete with the best emerging talent, the disruption to average writers is likely even greater.
We selected closed-source LLMs readily accessible to users without technical expertise: GPT-4o, Claude 3.5 Sonnet, and Gemini 1.5 Pro. We also tried open-weights Llama 3.1 model but its empirical performance was not good at following long context instructions at the time of our study. Both human experts and LLMs were given the same task: write an excerpt of up to 450 words emulating the style and voice of one of 50 internationally acclaimed authors, including Nobel laureates, Booker Prize winners, and Pulitzer Prize winners, spanning multiple continents and cultures. Among others, our study includes Nobel laureates Han Kang and Annie Ernaux; Booker Prize winners Salman Rushdie, Margaret Atwood, and George Saunders; and Pulitzer Prize winners Junot Díaz and Marilynne Robinson.
Our study examines shorter (up to 450-word) excerpts for two reasons. First, an excerpt of this length isolates the voice, style, and quality that writers must sustain across any longer work. Second, in its current form, an LLM cannot write long-form narratives that automatically balance coherence, quality, thematic consistency, and plot development across thousands of words. But it is merely a matter of time before LLMs reach that capacity. Even now, with human steering and iterative prompting, it is totally possible to produce long-form fiction and non-fiction using shorter model-generated excerpts. This is already happening in real life with startups such as Sudowrite or Genre Fiction publisher Inkitt  (?) specifically focusing on helping consumers write full books using AI. In self-publishing marketplaces such as Kindle, these AI-generated books have already gained popularity  (?, ?).
[…]
@@ 1 Results
Figure 2: (A-B) Forest plots showing odds ratios (OR) and 95% confidence intervals comparing AI and human experts in pairwise evaluations of stylistic fidelity (A) and writing quality (B) where values >>1 favor AI and values <<1 favor humans. Expert readers show preference for human writing when prompted in an in-context setting (OR = 0.16 and 0.13) but that changes when AI is fine-tuned (OR = 8.16 and 1.87). Lay readers have a harder time discriminating, given how they prefer the quality of AI writing even for in-context prompting (OR = 1.55). (C-D) Probability of choosing AI excerpts across individual language models for stylistic fidelity (C) and writing quality (D). Error bars represent 95% confidence intervals. Dashed line indicates chance performance (50%). (E) AI detection accuracy with chosen threshold of $\tau$=0.9 using two state-of-the-art AI detectors (Pangram and GPTZero). Human written text was never misclassified (0.00), in-context AI was detected with 97% accuracy by Pangram and 91% by GPTZero, but fine-tuned AI evaded detection 97% of the time (0.03) for Pangram and 100% of the time in case of GPTZero. (F) Relationship between AI detectability (Pangram) and preference for writing quality across detection score bins. For in-context prompting setup, higher detection scores correlated with lower AI preference (negative slope). This relationship disappeared when AI is fine-tuned (flat slopes). Fine-tuning on an author’s complete oeuvre gets rid of AI quirks while achieving expert-level performance. n = 28 expert readers, 131 lay readers; 3,840 pairwise comparisons with robust clustered standard errors.
@@ 1.1 Overall Performance Comparisons
Our final data consist of 3,840 paired-choice tasks, with judgments from 28 MFA experts and 131 lay readers. We fit a logit model for each outcome and condition and include dummies for writer type, reader group, and their interaction. We further employ CR2 cluster robust standard errors (?) clustered at the reader-level to account for within-reader correlation in ratings. Our hypotheses, outcomes, design, and analysis closely follow our OSF pre-registration (SI Sections S4-S8); minor deviations are detailed in SI Section S9.
Figure 2A–B presents the odds ratios, with corresponding predicted probabilities shown in Figure 2C–D. Under in-context prompting, expert readers demonstrated strong preference for human-written text. Odds ratios were 0.16 (95% CI: 0.08–0.29, $p<10^{-8}$) for stylistic fidelity and 0.13 (95% CI: 0.06–0.28, $p<10^{-7}$) for writing quality, indicating six- to eight-fold preferences for human excerpts (Fig. 2A–B). Lay readers showed no significant preference regarding stylistic fidelity (OR = 0.86, 95% CI: 0.62–1.19, $p=0.37$) but favored AI-generated text for writing quality (OR = 1.55, 95% CI: 1.09–2.21, $p=0.014$), selecting AI excerpts in 61% of writing quality trials (Fig. 2C–D). Inter-rater agreement reflected this divergence: expert readers achieved $\kappa=0.58$ for stylistic fidelity and $\kappa=0.41$ for writing quality, while lay readers showed minimal agreement among themselves ($\kappa=0.12$ and $\kappa=0.15$, respectively). Given that lay assessments of literary quality and style reflect inherently diverse tastes, we anticipated lower inter-rater agreement compared to expert evaluators. The writer-type $\times$ reader-type interaction was significant for both outcomes ($\chi^{2}_{(3)}=24.9$, $p=1.6\times 10^{-5}$ for fidelity; $\chi^{2}_{(3)}=37.6$, $p=3.5\times 10^{-8}$ for quality).
@@ 1.2 AI Detection and Stylometric Analysis
We probe whether differences in AI detectability can account for these preference reversals. Pangram, a state-of-the-art AI detection tool (?, ?, ?), correctly classified 97% of in-context prompted texts as machine-generated but only 3% of fine-tuned texts were classified as AI-generated (Fig. 2E).
GPTZero, another state-of-the-art AI detection, tool showed comparable performance but had higher false-positive-rate, so for our subsequent analyses we stuck to Pangram.
@@ 1.3 Author-Level Performance Heterogeneity
Figure 4: Fine-tuning substantially reduces stylometric signatures of AI text, improves stylistic fidelity and perceived writing quality over in-context prompting, and substantially cuts costs of producing first draft versus professional writers. (A) Mediation analysis linking stylometric features $\rightarrow$ AI-detector score $\rightarrow$ human preference. Standardized logistic coefficients with 95% CIs are shown for three features for in-context prompting (red) and fine-tuned models (green). Cliché density mediates 16.4% of the detector effect on choice for in-context prompting but only 1.3% for author fine-tuned models; all three features together mediate 25.4% vs -3.2% for in-context prompted and fine-tuned models respectively. (B) “Fine-tuning premium,” defined as $\Delta$ = P(prefer fine-tuned over human) - P(prefer in-context over human), as a function of fine-tuning corpus size. Top: stylistic fidelity; bottom: writing quality. Points are authors; colors denote improvement (green, $\Delta>0$), no change (gray), or degradation (red, $\Delta<0$). Median $\Delta$: +41.7 (fidelity) and +16.7 percentage points (quality). (C) Cost to produce 100,000 words of raw text vs. publishable prose. Expert writers in our study would earn \$25,000 for a 100k-word novel-length manuscript (red). By contrast, AI pipelines can generate 100k words of raw text for \$25–\$276 depending on fine-tuning corpus size (green bars = fine-tuning; hatched = in-context prompting, \$3). This figure reflects direct compute/API costs only, not the additional human steering, chunking, and editing required to turn raw AI text into a cohesive publishable work. Authors ordered by total AI cost.
@@ 1.4 Cost Analysis
The weak dependence of performance on fine-tuning corpus size has direct economic implications. Model fine-tuning costs ranged from \$25 to \$276 per author (median = \$81), assuming API-based fine-tuning at \$25 per million tokens plus \$3 for generating 100,000 words of raw text (Fig. 4C). These costs represent approximately 0.3% of what expert writers in our study would charge for a novel-length (100,000 words) manuscript. It should be noted that this comparison reflects raw generation costs before the human steering and editing required to transform AI outputs into publishable works. Despite this caveat, the minimal investment yielded outputs that achieved expert-level performance for most authors. Performance gains were uncorrelated with both corpus size and fine-tuning cost, indicating that computational scale did not drive improvements. The 99.7% reduction in raw generation costs, coupled with superior quality ratings for the majority of authors, underscores the potential for substantial producer-surplus shifts and market displacement.